# Paper/main_online_appendix

### v3

@@ Appendix D Appendix: Experiments
Here we provide more details on the design and pre-registration of our choice and valuation experiments experiments. Screenshots of experimental instructions, comprehension checks, and sample choice interfaces for these experiments are compiled in Online Appendix I.
@@ D.1 Multi-Attribute Binary Choice
The main sample consists of 80 two-attribute problems, 432 three-attribute problems, and 104 four-attribute problems. The three-attribute choice options are characterized by a monthly fee, a per-GB usage rate (where the fictional consumer has a monthly usage of 6 GB), and an annual device cost; the two-attribute choice options consist only of a monthly fee and usage rate, and the four-attribute choice options additionally contain a quarterly wi-fi charge. The two-attribute problems are generated by drawing a value difference (in bonus payment terms) from one of two values in {\$$\{\$3.84,\$5.76\}$ and an $L_{1}$-ratio from one of 12 values in $\{1.00,0.94,0.89,0.84,0.80,0.76,0.70,0.59,0.48,0.39,0.30,0.20\}$. [^46] Due to rounding in the attribute values, the actual $L_{1}$ ratios of the problems deviate slightly from these values. The three- and four-attribute problems are generated by similarly drawing a value difference and $L_{1}$ ratio value, which determines the summed attribute-wise advantages and disadvantages in the comparison, and randomizing how the advantage and disadvantages are split across the attributes.
@@ D.2 Intertemporal Binary Choice
In the broad problem sample, choice options contain either one or two payouts; in total, there are 300 1-payout vs. 1-payout choice problems, 300 1-payout vs. 2-payout choice problems, and 300 2-payout vs. 2-payout choice problems. For each choice problem, the options are generated by sampling payout amounts and payout delays. The delays of each payout (in days) are drawn from {0, 12, 24, 48, 72, 108, 144, 180, 216, 264, 312, 360, 420, 480, 540, 600, 660, 720}, and the monetary amount of each payout is drawn from {\$$\{\$0,\$0.50,...,\$20\}$ for two-payout options and {\$$\{\$0,\$0.50,...,\$40\}$ for one-payout options. Rather than uniformly sampling from these ranges, we employ a sampling procedure that 1) undersamples dominance problems, 2) excludes problems involving very large value differences and problems near indifference, and 3) stratifies by CPF ratio and value difference (computed using a benchmark discount factor).
@@ D.3.1 Lottery Preference Reversals: Experimental Details
Price List and Choice Construction. The experiment concerns 12 lottery options: 6 “base” options consisting of 3 high-risk and 3 low-risk options as described in Table 2, and 6 “scaled-up” options constructed by multiplying the payments in the base options by a scale factor of 1.6. The experiment consists of two parts, which subjects complete in random order: Binary choice and Valuation.
@@
In Valuation, subjects value all 12 options using one of two randomly assigned valuation modes: certainty equivalents and probability equivalents. In describing the price lists, we will denote each lottery being valued as (\$$(\$\lambda w,p)$, where $w$ is the base payoff amount, $p$ is the payoff probability, and $\lambda\in\{1,1.6\}$ is the scale factor. For certainty equivalent price lists, each lottery (\$$(\$\lambda w,p)$ is valued against a price list $Z=(z^{1},..,z^{n})$ of certain payments with evenly-spaced payoff differences, i.e. $z^{k}=(\$\lambda[w-(k-1)d),1)$ and $n=\lfloor w/d\rfloor$. We use $d=0.25$ for the low-risk lotteries and $d=1$ for the lottery (\$$(\$\lambda\cdot 19.5,0.23)$, and $d=1.25$ for the remaining high-risk lotteries. For probability equivalents, each lottery (\$$(\$\lambda w,p)$ is valued against a price list $Z=(z^{1},..,z^{n})$ of lotteries that pay off \$$\$\lambda\cdot 24$ with evenly-spaced payoff probability differences, i.e. $z^{k}=(\$\lambda\cdot 24,p-(k-1)d$) and $n=\lfloor p/d\rfloor$. We use $d=0.05$ for the low-risk lotteries and $d=0.01$ for the high-risk lotteries. For both valuation modes, the order of the valuation tasks is randomized subject to the constraint that no consecutive valuation tasks involve lotteries with the same payoff probability or scale factor.
@@ D.3.2 Intertemporal Preference Reversals: Experimental Details
Price List and Choice Construction. The experiment concerns 12 delayed payment options: 6 “base” options consisting of 3 high-delay and 3 low-delay options as described in Table 2, and 6 “scaled-up” options constructed by multiplying the payments in the base options by a scale factor of 1.6. The experiment consists of two parts, which subjects complete in random order: Binary choice and Valuation.
@@
In Valuation, subjects value all 12 options using one of two randomly assigned valuation modes: present value equivalents and time equivalents. In describing the price lists, we will denote each delayed payment being valued as $(\lambda m,t)$, where $m$ is the base payoff amount, $t$ is the payoff delay (in days), and $\lambda\in\{1,1.6\}$ is the scale factor. For present value equivalent price lists, each option (\$$(\$\lambda m,t)$ is valued against a price list $Z=(z^{1},..,z^{n})$ of immediate payments with evenly-spaced payoff differences, i.e. $z^{k}=(\$\lambda[w-(k-1)d),0)$ and $n=\lfloor w/d\rfloor$. We use $d=0.5$ for the low-delay options and $d=1.25$ for the high-delay options. For time equivalents, each option (\$$(\$\lambda m,t)$ is valued against a price list $Z=(z^{1},..,z^{n})$ of delayed payments $z^{k}=(\$\lambda\cdot 27.5,\tau_{k}$) and where $\tau_{k}=t+d_{k}$, for $\allowdisplaybreaks(d_{1},...,d_{n})=(0,7,15,30,45,60,90,120,180,240,300,360,420,480,540,600,660,720,840,960,1080)$. For both valuation modes, the order of the valuation tasks is randomized subject to the constraint that no consecutive valuation tasks involve lotteries with the same payoff delay or scale factor.
@@ D.3.3 Valuation Experiments: Experimental Details
Price List and Choice Construction. Subjects complete two parts of the experiment, in random order: Risk and Time. In Risk, subjects complete 12 multiple price list valuation tasks corresponding to one of two randomly selected valuation modes: certainty equivalents and probability equivalents. In Time, subjects complete 12 multiple price list valuation tasks corresponding to one of two randomly selected valuation modes: present value equivalents and time equivalents.
For certainty equivalents, subjects value a simple lottery $l=(\overline{w},p_{l})$, against a price list $Z=\{z^{1},...,z^{n}\}$ of certain payments adapted to $l$, with $n=19$. For probability equivalents, subjects value a certain payment $c=(w_{c},1)$ against a probability list $Z=\{z^{1},...,z^{n}\}$ of yardstick lotteries $z^{k}=(\overline{w},p_{k})$ adapted to $c$, with $n=21$. We draw $\overline{w}$ from {\$$\{\$9,\$18,\$27\}$. For certainty equivalents, we draw $p_{l}$ from $\{0.03,0.05,0.10,0.25,0.5,0.75,0.90,0.95,0.97\}$. For probability equivalents, we draw $w_{c}$ so that $w_{c}/\overline{w}\in\{0.033,0.056,0.11,0.25,0.5,0.75,0.89,0.944,0.967\}$. Subjects complete 12 price lists randomly selected from the 27 possible price lists; the order of the price lists are randomized, subject to the constraint that no consecutive price list contains the same payoff probability $p_{l}$ (normalized payment $w_{c}/\overline{w}$) for certainty equivalents (probability equivalents).
@@ Appendix E Appendix: Structural Specifications
$G(r)$ =$=\begin{dcases}(1-\kappa)-(0.5-\kappa)\frac{(1-r)^{\gamma}}{(r^{\psi}+(1-r)^{\psi})^{1/\psi}}&r\geq 0\\ \kappa+(0.5-\kappa)\frac{(1+r)^{\gamma}}{(r^{\psi}+(1-r)^{\psi})^{1/\psi}}&r<0\end{dcases}$
@@ E.2 Intertemporal Choice
Hyperbolic Discounting. We use the discount function proposed in Loewenstein and Prelec, (1992):
@@ E.3 Choice Under Risk
CDF Complexity. We estimate two versions of our model: one that assumes risk neutrality, and one that allows for utility curvature. In the risk neutral model, choice probabilities are given by
@@
In the model that allows for utility curvature, choice probabilities are given by
@@ Appendix F Appendix: Model Completeness and Restrictiveness
We consider prediction rules of the form $p:\mathcal{C}\to\mathcal{Q}$ and denote the set of such functions by $\overline{\mathcal{P}}$; $p(c)$ denotes the rate of choosing option $a$ in choice problem $c$ under prediction rule $p$. We are interested in evaluating the completeness and restrictiveness of a parametric model $\mathcal{P}_{\Theta}=\{p_{\theta}\}_{\theta\in\Theta}$. Completeness and restrictiveness are evaluated with respect to a base model $p^{b}$: we take $p^{b}$ to be the constant prediction rule $p^{b}(c)=\sum_{c\in\mathcal{C}}\frac{n_{c}}{\sum_{c^{\prime}\in\mathcal{C}}n_{c^{\prime}}}r(c)$ outputting the average choice rate in lottery and intertemporal choice, and to be the fitted Distortion-Free Logit model in multiattribute choice. [^47] In multiattribute choice, we use the Distortion-Free Logit model instead of the constant predictor as the base model since the constant predictor has lower prediction loss than Distortion-Free Logit; in the framework of Fudenberg et al., (2022, 2023), the base model should have higher prediction error than the model under evaluation.
@@ F.1 Completeness
Implementation. To form $p^{*}$, we use an ensemble approach to build a predictor that maps problem features into choice probabilities, where we combine parametric model predictions with those of a neural network trained on the data. This ensemble predictor is described in detail in Online Appendix G.
@@ F.2 Restrictiveness
Where $>$ denotes the domain-specific dominance notion. [^48] That is, $>$ denotes attribute-wise dominance, first-order-stochastic dominance, and temporal dominance in multiattribute, lottery, and intertemporal choice, respectively. Every model in Section E satisfies Weak Dominance, and every model satisfies Monotonicity except for the Salience model.
@@
Due to the high-dimensional nature of restrictions characterizing $\mathcal{P}$ – the Monotonicity restriction generates 648, 6616, and 93529 independent pairwise inequality constraints between choice rates in our datasets for multiattribute, intertemporal, and lottery choice, respectively – standard methods of sampling uniformly from $\mathcal{P}$, such as rejection sampling, are computationally infeasible. Instead, we approximate uniform draws from $\mathcal{P}$ by implementing a hit-and-run sampler (Smith,, 1984), a Markov chain Monte-Carlo algorithm designed to efficiently sample from a uniform distribution over a convex body. Online Appendix H describes this sampling procedure in full.

### v4

@@ Appendix D Appendix: Experiments
Here we provide more details on the design and pre-registration of our choice and valuation experiments experiments. Screenshots of experimental instructions, comprehension checks, and sample choice interfaces for these experiments are compiled in Supplemental Appendix I.
@@ D.1 Multi-Attribute Binary Choice
The main sample consists of 80 two-attribute problems, 432 three-attribute problems, and 104 four-attribute problems. The three-attribute choice options consist of a monthly fee, a per-GB usage rate (where the fictional consumer has a monthly usage of 6 GB), and an annual device cost; the two-attribute choice options consist of a monthly fee and usage rate, and the four-attribute choice options additionally contain a quarterly wi-fi charge. The two-attribute problems are generated by drawing a value difference (in bonus payment terms) from one of two values in $\{\$3.84,\$5.76\}$ and an $L_{1}$-ratio from one of 12 values in $\{1.00,0.94,0.89,0.84,0.80,0.76,0.70,0.59,0.48,0.39,0.30,0.20\}$. [^39] Due to rounding in the attribute values, the actual $L_{1}$ ratios deviate slightly from these values. The three- and four-attribute problems are generated by similarly drawing a value difference and $L_{1}$ ratio value, which determines the summed attribute-wise advantages and disadvantages in the comparison, and randomizing how the advantages and disadvantages are split across the attributes.
@@ D.2 Intertemporal Binary Choice
In the broad problem sample, choice options contain either one or two payouts; in total, there are 300 1-payout vs. 1-payout choice problems, 300 1-payout vs. 2-payout choice problems, and 300 2-payout vs. 2-payout choice problems. For each choice problem, the options are generated by sampling payout amounts and payout delays. The delays of each payout (in days) are drawn from {0, 12, 24, 48, 72, 108, 144, 180, 216, 264, 312, 360, 420, 480, 540, 600, 660, 720}, and the monetary amount of each payout is drawn from $\{\$0,\$0.50,...,\$20\}$ for two-payout options and $\{\$0,\$0.50,...,\$40\}$ for one-payout options. Rather than uniformly sampling from these ranges, we employ a sampling procedure that 1) undersamples dominance problems, 2) excludes problems involving very large value differences and problems near indifference, and 3) stratifies by CPF ratio and value difference (computed using a benchmark discount factor).
@@ D.3.1 Lottery Preference Reversals: Experimental Details
Price List and Choice Construction. The experiment concerns 12 lotteries: 6 “base” options consisting of 3 high-risk and 3 low-risk options as described in Table 2, and 6 “scaled-up” options constructed by multiplying the base option payments by a factor of 1.6. The experiment contains two parts, the order of which is randomized: Binary choice and Valuation.
@@
In Valuation, subjects value all 12 options using one of two randomly assigned valuation modes: certainty equivalents and probability equivalents. In describing the price lists, we will denote each lottery being valued as $(\$\lambda w,p)$, where $w$ is the base payoff amount, $p$ is the payoff probability, and $\lambda\in\{1,1.6\}$ is the scale factor. For certainty equivalent price lists, each lottery $(\$\lambda w,p)$ is valued against a price list $Z=(z^{1},..,z^{n})$ of certain payments with evenly-spaced payoff differences, i.e. $z^{k}=(\$\lambda[w-(k-1)d),1)$ and $n=\lfloor w/d\rfloor$. We use $d=0.25$ for the low-risk lotteries and $d=1$ for the lottery $(\$\lambda\cdot 19.5,0.23)$, and $d=1.25$ for the remaining high-risk lotteries. For probability equivalents, each lottery $(\$\lambda w,p)$ is valued against a price list $Z=(z^{1},..,z^{n})$ of lotteries that pay off $\lambda\cdot 24$ with evenly-spaced payoff probability differences, i.e. $z^{k}=(\$\lambda\cdot 24,p-(k-1)d$) and $n=\lfloor p/d\rfloor$. We use $d=0.05$ for the low-risk lotteries and $d=0.01$ for the high-risk lotteries. The order of the valuation tasks is randomized subject to the constraint that no consecutive valuation tasks involve lotteries with the same payoff probability or scale factor.
@@ D.3.2 Intertemporal Preference Reversals: Experimental Details
Price List and Choice Construction. The experiment concerns 12 delayed payments: 6 “base” options consisting of 3 high-delay and 3 low-delay options as described in Table 2, and 6 “scaled-up” options constructed by multiplying the base options payments by a scale factor of 1.6. The experiment consists of two parts, the order of which is randomized: Binary choice and Valuation.
@@
In Valuation, subjects value all 12 options using one of two randomly assigned valuation modes: present value equivalents and time equivalents. In describing the price lists, we will denote each delayed payment being valued as $(\lambda m,t)$, where $m$ is the base payoff amount, $t$ is the payoff delay (in days), and $\lambda\in\{1,1.6\}$ is the scale factor. For present value equivalent price lists, each option $(\$\lambda m,t)$ is valued against a price list $Z=(z^{1},..,z^{n})$ of immediate payments with evenly-spaced payoff differences, i.e. $z^{k}=(\$\lambda[w-(k-1)d),0)$ and $n=\lfloor w/d\rfloor$. We use $d=0.5$ for the low-delay options and $d=1.25$ for the high-delay options. For time equivalents, each option $(\$\lambda m,t)$ is valued against a price list $Z=(z^{1},..,z^{n})$ of delayed payments $z^{k}=(\$\lambda\cdot 27.5,\tau_{k}$) and where $\tau_{k}=t+d_{k}$, for $\allowdisplaybreaks(d_{1},...,d_{n})=(0,7,15,30,45,60,90,120,180,240,300,360,420,480,540,600,660,720,840,960,1080)$. The order of the valuation tasks is randomized subject to the constraint that no consecutive valuation tasks involve lotteries with the same payoff delay or scale factor.
@@ D.3.3 Valuation Experiments: Experimental Details
Price List and Choice Construction. Subjects complete two parts of the experiment, in random order: Risk and Time. In Risk (Time), subjects complete 12 multiple price list valuation tasks using one of two randomly selected valuation modes: certainty equivalents and probability equivalents (present value equivalents and time equivalents).
For certainty equivalents, subjects value a simple lottery $l=(\overline{w},p_{l})$, against a price list $Z=\{z^{1},...,z^{n}\}$ of certain payments adapted to $l$, with $n=19$. For probability equivalents, subjects value a certain payment $c=(w_{c},1)$ against a probability list $Z=\{z^{1},...,z^{n}\}$ of yardstick lotteries $z^{k}=(\overline{w},p_{k})$ adapted to $c$, with $n=21$. We draw $\overline{w}$ from $\{\$9,\$18,\$27\}$. For certainty equivalents, we draw $p_{l}$ from $\{0.03,0.05,0.10,0.25,0.5,0.75,0.90,0.95,0.97\}$. For probability equivalents, we draw $w_{c}$ so that $w_{c}/\overline{w}\in\{0.033,0.056,0.11,0.25,0.5,0.75,0.89,0.944,0.967\}$. Subjects complete 12 price lists randomly selected from the 27 possible price lists; the order of the price lists are randomized, subject to the constraint that no consecutive price list contains the same payoff probability $p_{l}$ (normalized payment $w_{c}/\overline{w}$) for certainty equivalents (probability equivalents).
@@ Appendix E Appendix: Structural Specifications
$G(r)$ $=\begin{dcases}(1-\kappa)-(0.5-\kappa)\frac{(1-r)^{\gamma}}{(r^{\psi}+(1-r)^{\psi})^{1/\psi}}&r\geq 0\\ \kappa+(0.5-\kappa)\frac{(1+r)^{\gamma}}{(r^{\psi}+(1-r)^{\psi})^{1/\psi}}&r<0\end{dcases}$
@@ E.2 Intertemporal Choice
Hyperbolic Discounting. We use the Loewenstein and Prelec, (1992) discount function:
@@ E.3 Choice Under Risk
Simplicity Theory. The DM has EU preferences, but pay penalize (or favor) lotteries with larger support. We follow Puri, (2025) in paramterizing the penalization term:
$\rho(x,y)$ $=\text{sgm}_{\eta}(V_{st}(x)-V_{st}(y))$
$V_{st}(x)$ $=EU_{sym}(x)+CA(|S_{x}|)$
$CA(s)$ $=\frac{\phi}{1+\exp(\upsilon(s-\mu))}-\frac{\phi}{1+\exp(\upsilon(1-\mu))}$
for $\alpha>0$. This model is parameterized by $(\eta,\alpha,\phi,\upsilon,\mu)$.
@@
CDF Complexity. We estimate two versions of our model: one that assumes risk neutrality, and one that allows for utility curvature. In the risk neutral model, we have
@@
In the model that allows for utility curvature, we have
@@ Appendix F Appendix: Model Completeness and Restrictiveness
We consider prediction rules of the form $p:\mathcal{C}\to\mathcal{Q}$ and denote the set of such functions by $\overline{\mathcal{P}}$; $p(c)$ denotes the rate of choosing option $a$ in choice problem $c$ under prediction rule $p$. We are interested in evaluating the completeness and restrictiveness of a parametric model $\mathcal{P}_{\Theta}=\{p_{\theta}\}_{\theta\in\Theta}$. Completeness and restrictiveness are evaluated with respect to a base model $p^{b}$: we take $p^{b}$ to be the constant prediction rule $p^{b}(c)=\sum_{c\in\mathcal{C}}\frac{n_{c}}{\sum_{c^{\prime}\in\mathcal{C}}n_{c^{\prime}}}r(c)$ outputting the average choice rate in lottery and intertemporal choice, and the fitted Distortion-Free Logit model in multiattribute choice. [^40] In multiattribute choice, we use the Distortion-Free Logit model instead of the constant predictor as the base model since the constant predictor has lower prediction loss than Distortion-Free Logit; in the framework of Fudenberg et al., (2022, 2023), the base model should have higher loss than the model under evaluation.
@@ F.1 Completeness
Implementation. To form $p^{*}$, we use an ensemble approach to build a predictor that maps problem features into choice probabilities, where we combine parametric model predictions with those of a neural network trained on the data. This ensemble predictor is described in detail in Supplemental Appendix G.
@@ F.2 Restrictiveness
Where >> denotes the domain-specific dominance notion. [^41] That is, >> denotes attribute-wise dominance, first-order-stochastic dominance, and temporal dominance in multiattribute, lottery, and intertemporal choice, respectively. Every model in Section E satisfies Weak Dominance except for Simplicity Theory, and every model satisfies Monotonicity except for Simplicity Theory and the Salience model.
@@
Due to the high-dimensional nature of restrictions characterizing $\mathcal{P}$ – the Monotonicity restriction generates 648, 6616, and 93529 independent pairwise inequality constraints between choice rates in our multiattribute, intertemporal, and lottery choice data, respectively – standard methods of sampling uniformly from $\mathcal{P}$ such as rejection sampling are computationally infeasible. Instead, we approximate uniform draws from $\mathcal{P}$ by implementing a hit-and-run sampler (Smith,, 1984), a Markov chain Monte-Carlo algorithm designed to efficiently sample from a uniform distribution over a convex body. Supplemental Appendix H describes this sampling procedure in full.